Pamela has a solution of 50% acid and another of 80% acid. How many ounces of each will make 300 ounces of a solution that is 72% acid?
Answer by Alan3354(69443)

Pamela has a solution of 50% acid and another of 80% acid. How many ounces of each will make 300 ounces of a solution that is 72% acid?
---------------
f = 50%
e = 80%
------
e + f = 300 (total solution)
80e + 50f = 72*300 (total acid)
----------
Solve for e & f